OpenStudy (anonymous):

HELP PLEASE?!? What is the density of a 35 g rock that displaces 7 ml of water??

OpenStudy (anonymous):

I need quick help please!

OpenStudy (ajprincess):

density=mass/volume

OpenStudy (anonymous):

ok, so what would it be??

OpenStudy (anonymous):

Would I just divide the 2 numbers?

OpenStudy (ajprincess):

7ml=7cm^3 d=35/7(g/cm^3) =5(g/cm^3)

OpenStudy (anonymous):

Ok, I get it now! Thanks

OpenStudy (aravindg):

density=mass/volume !!!!

OpenStudy (anonymous):

Thanks

OpenStudy (ajprincess):

yw.